OpenStudy (anonymous):

2. A chemist wishes to make 3 liters of a 7% acid solution by mixing a 9% acid solution and a 4% acid solution. How many liters of each solution should the chemist use? Write your answer as a complete sentence. Be sure to: • Define your variable and expressions for the quantities. • Write an equation that models the problem. • Solve the equation. • State the answer in a complete sentence

OpenStudy (anonymous):

Your two equations should look like this:\[x+y=3\]\[.09x+.04y=(.07)(3)\]

OpenStudy (anonymous):

@chucky00 what did you get on this, can you please tell me.

OpenStudy (anonymous):

x= the volume of 9% solution y= the volume of 4% solution You are given the total amount of solution, which is 3 liters. Since we are adding a volume of 9% solution, and a volume of 4% solution together to get a total solution of 3 liters. We have the first equation: x + y = 3

OpenStudy (anonymous):

Write an equation that models the problem. is x + y = 3

OpenStudy (anonymous):

The 4% solution is 4% per liter of volume. The 9% solution is 9% per liter of volume. Since we are making a solution of 7% per liter of volume and we want 3 liters, we change 7% to a decimal and multiply by 3. Change the 4% and 9% to decimals equivalents, .04 and .09. So our second equation is .09x +.04y = (.07)3)

OpenStudy (anonymous):

Solve the equation.?

OpenStudy (anonymous):

Rearrange the first equation. x+y=3 x=3-y Substitute into the second equation. .09(3-y) +.04y=.21 .27-.09y+.04y=.21 -.05y=-.06 y=1.2 liters Substitute y=1.2 into the first equation and solve for x x+1.2=3 x=1.8 liters Check both values in the second equation .09(1.8)+.04(1.2)=.21 .162+.048=.21 .21=.21 It checks

OpenStudy (anonymous):

x=(3−y) can it be that for Write an equation that models the problem.

OpenStudy (anonymous):

It models the first equation. But since you have two unknowns, you need two equations to solve the problem.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

can you make it shorter for Solve the equation.

OpenStudy (anonymous):

@marshallinwashington then I'm done

OpenStudy (anonymous):

You can use the elimination method to solve two equations. But its not that much shorter. You can go watch some videos on youtube that will show you how to solve two equations using the elimination method.

OpenStudy (anonymous):

like for example 4+8*2 4+16 20 like that step by step to make it way easier cause I was kinda confused

OpenStudy (anonymous):

@marshallinwashington

OpenStudy (anonymous):

the answer is x=1.8 liters

OpenStudy (anonymous):

@marshallinwashington then I'm done because I know you must be sick of me

OpenStudy (anonymous):

The answer is: The chemist uses 1.8 liters of 9% solution and 1.2 liters of 4% solution to make 3 liters of 7% solution.

OpenStudy (anonymous):

thanks sorry for bugging you

OpenStudy (anonymous):

That's OK. I wouldn't be here if I didn't want to help. I hope you got something out of it.

OpenStudy (anonymous):

i did I watch one of those im going to watch more

OpenStudy (anonymous):

If you do that, I predict great things for you. Hard work pays off.

OpenStudy (anonymous):

idk @needahighgradeA i never finished it i got busy XD

OpenStudy (anonymous):

o lol